OpenStudy (anonymous):

A recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function?

OpenStudy (anonymous):

@jhonyy9 do you understand this?

OpenStudy (anonymous):

Oops! A.) 5, 1, 7, 13, 19, ... B.) 5, -1, -7, -13, -19, ... C.) 5, 11, 17, 23, 29, ... D.) 5, -11, -17, -23, -29, ...

jhonyy9 (jhonyy9):

SO THEN F(1)=5 THAN ahhhhh Now is more understandably

jhonyy9 (jhonyy9):

try it for n=2 you get f(2) = f(2-1) -6 = f(1) -6 = 5-6=-1

OpenStudy (anonymous):

And you repeat that pattern, right? so the answer is B?

jhonyy9 (jhonyy9):

n=3 f(3) = f(3-1)-6 = f(2) -6 = -1-6=-7

jhonyy9 (jhonyy9):

yes, now hope you understand it sure right

OpenStudy (anonymous):

Yes, thank you!

jhonyy9 (jhonyy9):

was my pleasure bye

OpenStudy (anonymous):

Can you help me with another?